OpenStudy (anonymous):

solve the linear equation below. y = -2x + 19 y = X + 7

OpenStudy (anonymous):

X + 7 = -2x +19 2X+X+7=19 3X=12 X=12/3 x=4

OpenStudy (anonymous):

does that make sense to you ?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Well when u look at your equation there are two variables X and Y , we need to have one variable to solve the equation. So they have given you to get rid of the Y variable. I just replaced the Y variable with what the equation told me Y was equal to. so then I had only one variable to deal with. X :)

OpenStudy (anonymous):

take a few minutes to study it and you will understand